### v1

# Generalization Analysis for Bayesian Optimal Experiment Design under Model Misspecification

Abstract
In many settings in science and industry, such as drug discovery and clinical trials, a central challenge is designing experiments under time and budget constraints. Bayesian Optimal Experimental Design (BOED) is a paradigm to pick maximally informative designs that has been increasingly applied to such problems. During training, BOED selects inputs according to a pre-determined acquisition criterion. During testing, the model learned during training encounters a naturally occurring distribution of test samples. This leads to an instance of covariate shift, where the train and test samples are drawn from different distributions. Prior work has shown that in the presence of model misspecification, covariate shift amplifies generalization error. Our first contribution is to provide a mathematical decomposition of generalization error that reveals key contributors to generalization error in the presence of model misspecification. We show that generalization error under misspecification is the result of, in addition to covariate shift, a phenomenon we term error (de-)amplification which has not been identified or studied in prior work. Our second contribution is to provide a detailed empirical analysis to show that methods that result in representative and de-amplifying training data increase generalization performance. Our third contribution is to develop a novel acquisition function that mitigates the effects of model misspecification by including a term for representativeness and implicitly inducing de-amplification. Our experimental results demonstrate that our method outperforms traditional BOED in the presence of misspecification.
1 Introduction
Bayesian modeling is a principled approach to making inferences when data is scarce or costly. Most Bayesian machine learning methods are developed under the assumption that the true data-generating process (DGP) is included in the chosen model family [bernardo2009bayesian]. However, in complex real-world environments, this assumption rarely holds and the learner has no access to the true DGP [uppal2003model, sloman2024bayesian]. As a result, the true DGP often lies outside of the assumed model family. This inevitability of the phenomenon of model misspecification [walker2013bayesian] is captured by the somewhat infamous saying that “all models are wrong” [George1976Science, box1980sampling]. Common causes of model misspecification include omitted variables [wooldridge2010econometric], mistaken beliefs about the structure of the error term (e.g., a failure to account for heteroskedasticity or autocorrelation) [greene2003econometric, grunwald2017inconsistency], or the choice of a misinformed or underexpressive model class [wooldridge2010econometric, dubova2025ockhams]. The consequences of model misspecification range from biased inferences [greene2003econometric, muller2013risk, caprio2023credal, bonhomme2022minimizing], unreliable approximations (e.g., in simulation-based inference methods [frazier2020model, lintusaari2017fundamentals, huang2023learning]), to suboptimal decision-making [sutton1998reinforcement, rainforth2024modern].
There is a substantial literature on the effects of model misspecification on Bayesian inference when data is independently and identically distributed (i.i.d.), or “passively” collected from the distribution on which the learner wants their inferences to generalize [kleijn2006misspecification, kleijn2012bernsteinvonmises, knoblauch2022optimizationcentric, walker2013bayesian, nott2023bayesian, kelly2025simulation]. However, in part because of the widespread availability of large datasets, active learning methods have become much more prevalently used [settles2009active]. These methods select the training data to tailor it to a specified learning objective [silberman1996active, farquhar2021statistical]. Active learning methods rely on the specified model twice: once to make inferences to fit training data, and again to select the data [konyushkova2017learning]. Model misspecification thus has a double impact on these methods, introducing potential bias in both the acquisition function and the resulting inferences. In particular, in the context of active learning, model misspecification can produce poor quality datasets [sugiyama2005active, bach2006active, ali2014active, vincent2017darc, farquhar2021statistical]. Understanding the consequences of model misspecification is of paramount importance to developing robust active learning methods.
In a Bayesian setting, Bayesian Optimal Experimental Design (BOED) is a natural and frequently used active learning method [rainforth2024modern]. BOED selects the optimal design by maximizing an acquisition function known as the expected information gain, enabling budget and time efficiency [rainforth2024modern, chaloner1995bayesian] in many applications, such as drug discovery [park2013bayesian], clinical trial design [chaloner1995bayesian], chemistry [walker2019bayesian, hickman2022bayesian], biology [kreutz2009systems, thompson2023integrating], and psychology [cavagnaro2010adaptive, myung2013tutorial]. While the limitations of BOED in the presence of model misspecification have been acknowledged in the literature, only a limited number of papers have investigated this limitation [rainforth2024modern, foster2021deep, sloman2022characterizing, schmitt2023detecting, ivanova2024step].
We provide a novel theoretical analysis of generalization error in the presence of model misspecification. Our analysis reveals that training datasets that lead to robustness to model misspecification have two properties: They are representative of the target data-generating distribution, and they are de-amplifying. The expected information gain does not include a term for either of these, and standard BOED can lead to training datasets that have neither of these characteristics. In this sense, standard BOED is not robust to model misspecification.
Unrepresentative training data. BOED selects samples to achieve a particular objective, and these samples likely do not reflect the distribution to which the learner would like to generalize. In other words, BOED induces a form of distribution shift, whereby the distribution used for (active) learning is different than the distribution used for evaluation. Recent work on the interaction between model misspecification and distribution shift has introduced the concept of misspecification amplification [amortila2024mitigating], whereby the generalization error attributable to misspecification is “amplified” by the density ratio between the test and training input distributions. A similar phenomenon has been observed in the context of BOED: In the presence of model misspecification, the generalization error in some settings has been shown to depend on both the degree of model misspecification and the extent of distribution shift [sloman2022characterizing].
De-amplifying training data. As our novel decomposition of generalization error shows, generalization performance depends not only the representativeness of the training data, but also on the way it interacts with model (mis)specification: Generalization performance is enhanced when training data is most representative of the test distribution, and also in regions where the model is most misspecified in a helpful (i.e., error-reducing) direction. We refer to this property as error “de-amplification” to stress that the effect is to counteract, rather than amplify, the effect of the misspecification.
Contributions. In this work, we explore the problem of BOED under model misspecification and make the following contributions:
  • Theoretical decomposition of generalization error. Prior work has primarily explored the effects of misspecification and distribution shift, overlooking the role of de-amplifying designs. We formally decompose generalization error into three components: (1) misspecification bias, (2) estimation bias, and (3) a novel term we introduce, error (de-)amplification. We also derive an upper bound on generalization error that characterizes its dependence on the representativeness of the training data, the degree to which these data are de-amplifying, and model misspecification.
  • Empirical analysis of BOED under misspecification. We evaluate both myopic and non-myopic BOED methods in terms of their robustness to model misspecification. Our results demonstrate that differences in performance can be accounted for by the terms in our decomposition. We provide a detailed empirical analysis supporting this effect: Methods that result in representative and de-amplifying training data increase generalization performance.
  • Novel acquisition function via representativeness and de-amplification. We propose a novel acquisition function designed to mitigate the effects of model misspecification by identifying designs that not only are informative, but are additionally representative and implicitly de-amplifying. Our empirical results show that the new acquisition outperforms traditional BOED in the presence of misspecification.
2 Preliminaries
2.1 Problem Setting
General formulation
A modeler aims to predict an observed variable $y\in\mathbb{R}^{m}$ which depends on a fully observable input (design) $\xi\in\mathbb{R}^{d}$. The relationship between the observed variable $y$ and the input $\xi$ is governed by a conditional distribution $y|\xi\sim P^{\star}$, referred to as the true data-generating process (DGP), which depends on the output of a true (possibly unknown) regression function $f^{\star}(\xi)$ and observation noise. To approximate the true DGP, the modeler proposes a hypothetical model $f(\xi,\theta),$ where $\theta\in\Theta$ represents the parameters within the fixed parameter space $\Theta$. The model class is denoted by $\mathcal{F}(\xi,\Theta)=\{f(\xi,\theta):\theta\in\Theta\}$. Model misspecification arises when the assumed model class $\mathcal{F}(\xi,\Theta)$ fails to capture the true DGP [ivanova2024step, sloman2022characterizing].
Definition 2.1 (Model misspecification).
Model misspecification occurs when the assumed model class $\mathcal{F}(\xi,\Theta)=\{f(\xi,\theta):\theta\in\Theta\}$ cannot mimic the true output $f^{\star}(\xi)$ for any parameter $\theta\in\Theta$. That is, the model is misspecified if
$f^{\star}(\xi)\notin\mathcal{F}(\xi,\Theta).$ (1)
In Bayesian inference, the modeler additionally specifies a prior distribution over the model parameters. According to this prior, the probability the data the learner will encounter is generated by a value $\theta$ is $p(\theta)$. Model fitting is carried out by updating the prior distribution using Bayes’ rule. The result is a posterior distribution which assigns to a $\theta$ a probability $p(\theta\mid y,\xi)\propto p(\theta)p(y\mid\theta,\xi)$. This process depends on both the assumed prior and the specified likelihood model. When the model is misspecified, i.e., the likelihood does not reflect the true DGP, the updated posterior becomes unreliable or biased [frazier2023reliable, oberauervariance] .
2.2 Bayesian Optimal Experimental Design
Bayesian Optimal Experimental Design (BOED) is a model-based framework to select the optimal design by maximizing the expected information about the parameter, enabling budget and time efficiency [rainforth2024modern, chaloner1995bayesian]. The primary objective is to find a design $\xi$ that is optimal in the sense that it provides the most information about the parameter $\theta$. The expected information gain (EIG) captures the amount of information gained about $\theta$ [dong2024variational, lindley1956measure]:
$\operatorname{EIG}(\xi)=\mathbb{E}_{p(y\mid\xi)}[\operatorname{IG}_{\theta}(\xi,y))]=\mathbb{E}_{p(\theta,y\mid\xi)}[\log p(y\mid\theta,\xi)-\log p(y\mid\xi)]$ (2)
The optimal design $\xi^{\star}$ is the design in a set of candidate designs $\Xi$ that maximizes the EIG:
$\xi^{\star}={\operatorname{argmax}_{\xi\in\Xi}}\operatorname{EIG}(\xi).$ (3)
Traditional BOED methods [foster2019variational, sebastiani1997bayesian], also called Bayesian Adaptive Design (BAD), iterate between making design decisions by evaluating Equation 3, and updating the underlying model through Bayesian inference to condition on data obtained so far. Traditional BOED is computationally expensive, due to substantial costs required to estimate and optimize $\operatorname{EIG}(\xi)$ and update the model at each step. To mitigate these limitations, policy-based algorithms [ivanova2024step, blau2022optimizing], such as deep adaptive design (DAD) [foster2021deep], which trains a neural network based on a set of simulated experimental trajectories, have been proposed. By fully amortizing the design selection process, policy-based methods eliminate the need for step-by-step online model updates, significantly reducing computational overhead.
2.3 Distribution Shift
Distribution shift is a well-known challenge in machine learning. It refers to the setting where the input (design) distribution differs between training and test phases. In BOED, training designs are selected via an acquisition criterion, while the model’s performance at test-time is evaluated on a given test distribution of interest. This mismatch induces a specific form of distribution shift known as covariate shift, where the distribution of inputs shifts (i.e., $p_{\text{train}}(\xi)\neq p_{\text{test}}(\xi)$) while the conditional output distribution remains unchanged (i.e., $p(y\mid\xi_{\text{train}})=p(y\mid\xi_{\text{test}})$). Prior work has studied the covariate shift induced by BOED [sugiyama2005active, ali2014active, sloman2022characterizing]. To address the potential resultant biases, density ratio estimation — whereby the training data are reweighted according to the estimated ratio between test and training input distributions — has proven effective [ge2023maximum].
3 Theoretical results
3.1 Decomposition of generalization error
Recent work has demonstrated that generalization error depends on an interaction between the degree of covariate shift (the degree to which the training data are unrepresentative of the test distribution) and of model misspecification [amortila2024mitigating, ge2023maximum, wen2014robust]. In this section, we show that generalization error additionally depends on the degree of presence of a phenomenon we term error (de-)amplification. We show that generalization error can be decomposed into three terms, reflecting separate contributions of the degree of misspecification bias, of estimation bias, and of error (de-)amplification.
Definition 3.1 (Generalization error ($R_{\mathrm{test}})$).
Let $\{(\xi_{i},y_{i})\}_{i=1}^{n}$ be a dataset of $n$ i.i.d. samples drawn from the true DGP $P^{\star}$. Denote by $f^{\star}(\xi)$ the true data-generating function. Let $\hat{f}^{(n)}(\xi)$ be a learned predictor, depending on training data $\{\xi_{1},\ldots{},\xi_{n}\}$. Let $d_{\text{test}}$ be the test data distribution. The generalization error is defined as
$R_{\mathrm{test}}(\hat{f}^{(n)}):=\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\hat{f}^{(n)}(\xi)-f^{\star}(\xi))^{2}\right].$ (4)
Proposition 3.2 (Generalization Error Decomposition).
Let $\bar{f}\in\mathcal{F}$ be the predictor that best approximates the true data-generating function $f^{\star}$, i.e.$\bar{f}=\arg\min_{f\in\mathcal{F}}R_{\mathrm{test}}(f)$. Equation 4 can be decomposed into the following equation. A detailed derivation is provided in Appendix A.
$R_{\mathrm{test}}(\hat{f}^{(n)})$ $=\underbrace{\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\bar{f}(\xi)-f^{\star}(\xi))^{2}\right]}_{\text{Misspecification Bias ($\mathbb{B}$)}}+\underbrace{\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))^{2}\right]}_{\text{Estimation Bias ($\mathbb{C}$)}}$ (5)
$\quad+\underbrace{2\,\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\bar{f}(\xi)-f^{\star}(\xi))(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))\right]}_{\text{Error (de-)amplification ($\widehat{\mathbb{A}}$)}}.$
In the well-specified case where the true function lies within the model class, $\bar{f}(\xi)=f^{\star}(\xi)$. In this case, both the bias and interaction terms vanish, and the generalization error reduces to:
$R_{\mathrm{test}}(\hat{f}^{(n)})=\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))^{2}\right],$ (6)
where the only quantity that depends on the training sample is $\hat{f}^{(n)}(\xi)$, since the population function $d_{\text{test}}$ and the best predictor $\bar{f}$ are fixed.
In the misspecified case where the true function lies outside the model class, $\bar{f}(\xi)\neq f^{\star}(\xi)$. In this case, all three terms in Equation 5 contribute to generalization error. The terms have the following interpretations:
  • Misspecification bias ($\mathbb{B}$) captures the discrepancy between the best predictor and the true data-generating function, and reflects the degree of model misspecification. This term is fixed and unaffected by the training data.
  • Estimation bias ($\mathbb{C}$) captures the discrepancy between the best predictor $\bar{f}\in\mathcal{F}$ and the predictor arrived at on the basis of finite training data. In the BOED setting, this training data depends on the modeler’s sequential evaluations of the EIG.
  • Error (de-)amplification ($\widehat{\mathbb{A}}$) measures the correlation between the extent and direction of the model’s bias and the extent and direction of the estimation error over the test distribution. This term can either amplify or mitigate the overall generalization error:
    • A positive correlation indicates that the directions of misspecification and estimation bias tend to coincide, which amplifies (increases) the generalization error.
    • A negative correlation indicates that the directions of misspecification and estimation bias tend to counteract each other, which de-amplifies (decreases) the generalization error.
3.2 An upper bound on generalization error with error (de-)amplification
While Proposition 3.2 provides valuable insights into the various contributors to generalization error, computing these terms requires evaluating the outputs of $f^{\star}$ and $\bar{f}$ in expectation over the test samples.
Of course, this is infeasible in practice. To understand and control generalization error during training, the learner requires a formulation that explicitly relates it to quantities available during training. Theorem 3.7 provides such a formulation. Theorem 3.7 shows that the learner can control generalization error by selecting designs that (i) are representative of the test distribution (reduce the degree of covariate shift), an insight we apply in our development of a novel acquisition function in Section 4, and (ii) are de-amplifying (have the potential to counteract the misspecification bias).
Theorem 3.7 builds on a result from [amortila2024mitigating]. In particular, we tighten the upper bound introduced by [amortila2024mitigating] to depend on the degree of (de-)amplification of the training samples. In their analysis, [amortila2024mitigating] introduce the following quantities:
  • Degree of covariate shift ($\mathbb{C}_{\infty}$) measures the worst-case distance between the selected training samples and the test samples.
  • Degree of misspecification ($\mathbb{B}_{\infty}$) measures the worst-case discrepancy between the DGP and the best predictor in the model class.
Their analysis requires the following assumptions:
Assumption 3.3 (The degree of covariate shift is bounded [amortila2024mitigating]).
We assume that $d_{\text{train}}(\xi)>0$ for all $\xi\in\Xi$ and the density ratio between the test and training input distributions is uniformly bounded [amortila2024mitigating, sugiyama2007direct]:
$\mathbb{C}_{\infty}:=\sup_{\xi\in{\Xi}}|\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}|<\infty$ (7)
where $d_{\text{train}}$ and $d_{\text{test}}$ represent the train and test distribution respectively, $\mathbb{C}_{\infty}\geq 1$ always.
Remark 3.4.
Assumption 3.3 requires that $\mathrm{supp}(d_{\text{test}})\subseteq\mathrm{supp}(d_{\text{train}})$. Often in the BOED setting, training samples are selected from the test distribution, and so $\mathrm{supp}(d_{\text{train}})\subseteq\mathrm{supp}(d_{\text{test}})$. Although in such cases $\mathbb{C}_{\infty}$ may not be an applicable measure of covariate shift, our results nevertheless provide insight about the connection between covariate shift and generalization error.
The assumption of a bounded density ratio is standard in the covariate shift literature and enables tractable analysis of distribution mismatch [amortila2024mitigating].
Assumption 3.5 (The degree of misspecification is bounded [amortila2024mitigating]).
Let $\mathcal{D}$ be a discrepancy measure between functions. Let $\mathcal{F}$ denote the model class, and let $\bar{f}\in\mathcal{F}$ be the predictor that best approximates the true data-generating function $f^{\star}$: $\bar{f}=\arg\min_{f\in\mathcal{F}}R_{\mathrm{test}}(f)$. We assume that the discrepancy between the predictive distribution induced by $\bar{f}$ and that of the true data-generating function $f^{\star}$ is bounded:
$\mathcal{D}(\bar{f},f^{\star}):=\|\bar{f}-f^{\star}\|_{\infty}=\sup_{\xi\in\Xi}\mid\bar{f}(\xi)-f^{\star}(\xi)\mid\leq\mathbb{B}_{\infty}$ (8)
where $\mathbb{B}_{\infty}\geq 0$ is referred to as the misspecification margin.
Notice that if $\mathbb{B}_{\infty}=0$, the model is well-specified, i.e., $f^{\star}\in\mathcal{F}$. On the other hand, if $\mathbb{B}_{\infty}>0$, the model is misspecified (i.e., $f^{\star}\notin\mathcal{F}$) and $\mathbb{B}_{\infty}$ quantifies the degree of misspecification.
Assumption 3.6 (Boundedness of function [amortila2024mitigating]).
For all $\xi\in\Xi$,
$\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq y_{\infty}\text{, }\|f^{\star}\|_{\infty}\leq y_{\infty}\text{, and }|y|\leq y_{\infty}$
for some $y_{\infty}>0$ and where $\|f\|_{\infty}=\sup_{\xi\in\Xi}\mid f(\xi)\mid$.
Our result.
Theorem 3.7 extends the result of [amortila2024mitigating] by explicitly characterizing the behavior of generalization error in a way that accounts for error (de-)amplification. This additional term is important to be considered within the limited training sample settings under BOED, as it captures how the interaction term between training data and the model misspecification can either amplify or de-amplify generalization error.
Theorem 3.7 (Generalization Error Bound under Covariate Shift with Amplification).
Let $\mathcal{F}$ be a finite model class, and let $f^{\star}$ denote the true regression function. Let $\hat{f}^{(n)}$ be the empirical risk minimizer over training data drawn from the distribution $d_{\mathrm{train}}$. Then, with probability at least $1-\delta$, the generalization error under covariate shift satisfies:
$R_{\text{test}}(\hat{f}^{(n)})\leq\mathbb{C}_{\infty}\cdot\left\{\begin{aligned} &\mathbb{B}_{\infty}^{2}+\frac{224y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-2\mathbb{A}(\hat{f}^{(n)}),&\text{if }\mathbb{A}(\hat{f}^{(n)})<0\\ &\mathbb{B}_{\infty}^{2}+\frac{128y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-\sqrt{3}\mathbb{A}(\hat{f}^{(n)}),&\text{if }0\leq\mathbb{A}(\hat{f}^{(n)})\end{aligned}\right.$ (9)
where $\mathbb{C}_{\infty}:=\sup_{\xi\in\Xi}\left|\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\right|$ and $\mathbb{A}(\hat{f}^{(n)})=\mathbb{E}_{\xi\sim d_{\text{train}}}\left[(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$. The proof can be found in Appendix B.
Remark 3.8 (Connection to Proposition 3.2).
This bound is consistent with the full generalization error decomposition structure, including the cross-term $\mathbb{A}(f)$, which captures the interaction between properties of the model (misspecification bias) and of the sampling strategy (estimation bias). Proposition 3.2 does not explicitly reveal how the representativeness and (de-)amplifying properties of the training data interact with model bias. In contrast, Theorem 3.7 makes this interaction explicit, providing a more interpretable perspective for understanding and controlling the generalization error during the training phase.
Remark 3.9.
Proposition 3.2 defines the generalization error based on the learned function $\hat{f}^{(n)}$ and the true data-generating function , $f^{\star}$ over the test distribution. However, since the true data-generating function $f^{\star}$ is unknown and outside the model class, it cannot offer guidance for decision-making. In contrast, although the $\mathbb{A}$ term that appears in Theorem 3.7 depends on the unknown $\bar{f}$, $\bar{f}$ is located in the model class $\mathcal{F}$. While $\bar{f}$ is inaccessible to the modeler, its location within $\mathcal{F}$ suggests that it can be better approximated in practice.
Remark 3.10.
Theorem 3.7 requires the assumption that $\hat{f}^{(n)}$ is the member of $\mathcal{F}$ that minimizes risk in the training data. Although our experiments adopt the Bayesian learning framework, Theorem 3.7 still provides valuable insights about the role of representative and (de-)amplifying training samples.
Theorem 3.7 reveals that the following factors contribute to the generalization error:
  • Representativeness of the training data ($\mathbb{C}_{\infty}$). The multiplicative factor of $\mathbb{C}_{\infty}$ implies that, by reducing $\mathbb{C}_{\infty}$, representative training data can reduce generalization error.
  • Misspecification bias ($\mathbb{B}_{\infty}$). Misspecification bias cannot be reduced by the training data. However, because of the multiplicative effect of $\mathbb{C}_{\infty}$, the effect of model misspecification on generalization error can be amplified by unrepresentative training samples [amortila2024mitigating].
  • Error (de-)amplification ($\mathbb{A}$). While this term is not computable in practice, as $\bar{f}$ is unknown to the learner, it captures a key component of generalization error: the interaction between the learner’s misspecification and estimation errors. Because it is not computable in practice, it cannot (unlike covariate shift) be explicitly represented in an acquisition function. However, we show empirically in Section 5 that some acquisition functions induce more error de-amplification than others, and in this sense are “implicitly” de-amplifying.
    The term $\mathbb{A}(\hat{f}^{(n)})$ can be interpreted as an average, across the training samples, of the (signed) estimation errors weighted by the (signed) misspecification errors. When $\mathbb{A}(\hat{f}^{(n)})\geq 0$ (error de-amplification), the learner’s misspecification and estimation biases tend to agree, resulting in de-amplifying behavior. In contrast, when $\mathbb{A}(\hat{f}^{(n)})<0$ (error amplification), the estimated model tends to misalign with the direction of the model misspecification, resulting in amplifying behaviour.
    One of the learner’s objectives is to reduce error amplification in the cases where it is present ($\mathbb{A}(\hat{f}^{(n)})<0$). Since misspecification bias is irreducible, this implies reducing the estimation error on designs that (at present) induce amplification. This can be seen by noticing that according to Theorem 3.7 generalization error depends negatively on $\mathbb{A}(\hat{f}^{(n)})$, i.e., on the amplification term evaluated at the training data samples. In other words, de-amplifying samples (samples that will produce training data that will reduce amplification) have a high corresponding value of $\mathbb{A}(\hat{f}^{(n)})$.
4 A novel acquisition function
Leveraging the insights from Theorem 3.7, we design an acquisition function that identifies designs that are representative and de-amplifying, while providing information about the parameter of interest. Although the acquisition function does not include an explicit term corresponding to de-amplification (since this property is not computable in practice; see Remark 3.9), our empirical results show that it nevertheless tends to select more de-amplifying samples than other BOED methods. To measure the distance between the distributions implied by two datasets, we use the maximum mean discrepancy(MMD).
To account for covariate shift, we modify the standard EIG acquisition function by introducing an MMD-based correction term. The idea is to encourage the selection of design points that not only have high information gain but also help reduce the difference between the distributions of training and test points. Specifically, we use the following form:
[…]
where $h_{t-1}$ is the history of selected designs before the time step $t$. This approach selects more representative samples than standard BOED, which is intended to improve generalization under model misspecification. This robust acquisition penalizes designs that are only representative or only informative; in other words, the design under the robust acquisition is both representative and informative. The hyperparameter $\lambda$ controls the tradeoff between informativeness and representativeness. When $\lambda$ tends to zero, the selected designs are informative, and our new acquisition function converges to the traditional BOED method.
5 Experiments
This section contains comparative experiments and analysis to explore which algorithm performs best in the presence of the model misspecification in two experimental paradigms: a toy example and a source location paradigm. We also empirically validate the theoretical results in Section 3.
We compare the following methods: A Random strategy selects designs from the test distribution at random. Bayesian adaptive design (BAD) [foster2019variational] selects designs according to the traditional BOED strategy, i.e., sequentially on each step of the experiment. Deep adaptive design (DAD) [foster2021deep] selects designs according to a policy-based BOED algorithm, i.e., fully amortizes the design process. Adjusted BAD (BAD-Adj.) selects designs according to our novel acquisition function.
Besides the relative performance of our novel acquisition function, we are also interested in how model misspecification affects the BAD vs. DAD algorithms. The iterative paradigm of BAD can directly interact with the true DGP at each iteration, potentially providing it with the opportunity to adapt when the model is misspecified. In contrast, DAD is trained offline and cannot adapt through interaction with the true DGP.
5.1 Toy example
Polynomial regression model The DGP is a degree-two polynomial regression model, $y=1+2x-0.5x^{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,0.1)$. In the misspecified case, we use a linear model to fit the data this model generates. In the well-specified case, we use a quadratic model to fit the data this model generates. Additional results with a variety of hyperparameter settings are given in Appendix C.1. Figure 1 presents the results for both well-specified and misspecified models across 20 runs. For each run, the design is adaptively selected, and the total number of designs is $T=10$.
Figure 1: Comparison of different design strategies under well-specified and misspecified models in both polynomial regression and source location experiments. Top row: MMD distance across methods; higher values indicate a greater degree of covariate shift. Bottom row: Generalization error across methods. Left column: Comparison of existing methods (Random, BAD, DAD) in the polynomial regression experiments. Middle column: Comparison of baseline methods (Random, BAD) and our proposed BAD-Adj with varying $\lambda$ in the polynomial regression experiments. Right column: Comparison of baseline methods (Random, BAD) and our proposed BAD-Adj with varying $\lambda$ in the source location experiments.
Relationship between representativeness and generalization error Figure 1 shows that, in the well-specified case, all methods yield similar generalization error (Figure 1(d)), regardless of the degree of covariate shift (Figure 1(a)). This indicates that covariate shift does not significantly impact generalization performance when the model is well-specified. Under model misspecification, however, covariate shift negatively affects model performance. As shown in Figure 1(a), the degree of covariate shift induced by BAD is higher than the degree of covariate shift induced by Random (Figure 1(a)). This higher degree of covariate shift translates into larger generalization errors (Figure 1(d)). This suggests that in the presence of model misspecification, unrepresentative designs (larger covariate shift) decrease model performance. Figures 1(b) and 1(e) show the performance of our acquisition function under various values of $\lambda$. For larger values of $\lambda$, we expect the representativeness term to dominate the acquisition function, resulting in a design distribution that resembles the test distribution. Figure 1(b) shows that when designs are more representative, generalization error is reduced (Figure 1(e)), consistent with the theoretical prediction introduced in Theorem 3.7. These results demonstrate again that representative designs effectively reduce estimation bias and improve generalization performance.
Capacity to reduce error amplification To illustrate the behavior of each term in Proposition 3.2 at each design, we pick one of the 20 runs, compute each term in the error decomposition after 10 experimental steps, and plot the values across the design range [-4,4]. The results can be found in Figure 2.
This figure also identifies whether a training design is de-amplifying or amplifying. Recall from Section 3.2 that de-amplifying designs have a high corresponding value of $\mathbb{A}(\hat{f}^{(n)})$ (shown by the blue dotted lines in Figure 2). A training design is de-amplifying when the blue line is above zero at the value of the design. Conversely, the design is amplifying when the blue curve is below zero at the value of the design. Intuitively, from Theorem 3.7, under a limited design budget, the next design should be chosen where the interaction term is largest—that is, to reduce estimation error where the blue curve is highest—as this de-amplifies error arising from misspecification bias.
In the training phase, the Random strategy selects representative designs, but does not include a term to target (de-)amplification. As shown in Figure 2(a), this strategy sometimes selects de-amplifying designs (a positive value of the blue line) by chance, resulting in an error (de-)amplification term whose expectation across all training designs is slightly negative. In contrast, the acquisition functions used by both BAD and DAD lack explicit terms corresponding to either (de-)amplification or representativeness. Figures 2(b) and 2(c) suggest that the designs selected by these methods nevertheless lead to a generally negative amplification term at the training data. Even though DAD picks more representative designs compared to BAD, in general, the designs selected by DAD are more amplifying (the corresponding value of the blue line is further below zero) than those selected by BAD. Consistent with the result in Theorem 3.7 and as shown in Figure 1(a), this leads DAD to produce higher generalization error on the test distribution. Figure 2(d) shows that our proposed method, BAD-Adj, produces a training set that includes both amplifying and de-amplifying designs, suggesting that even though our proposed acquisition function does not target de-amplifying regions directly, it still has a higher chance of selecting de-amplifying designs than BAD or DAD. Figure 2(d) also shows that the overall degree of (de-)amplification induced by the designs selected by BAD-Adj is close to zero. These results suggest that, like Random, BAD-Adj sometimes selects de-amplifying designs by chance. These results suggest that the selection of de-amplifying designs can help reduce generalization error (as shown in Figure 1(e)), consistent with the theoretical result established in Theorem 3.7.
Figure 2: Comparison of different design strategies (Random, BAD, DAD, and proposed BAD-Adj) under the misspecified model in the polynomial regression experimental settings. The curves represent the local contribution of each term at different values of $\xi$. The green, orange, and blue lines represent misspecification term $(\bar{f}(\xi)-f^{*}(\xi))$, estimation term $(f(\xi)-\bar{f}(\xi))$, and error amplification (interaction) term $(f(\xi)-\bar{f}(\xi))(\bar{f}(\xi)-f^{*}(\xi))$, respectively. The blue dotted lines indicate the designs selected from iterations 1–10, with blue labels showing the corresponding iteration numbers.
5.2 Source localization experiments
The acoustic energy attenuation model simulates the total intensity at location $\xi$ from multiple sources at locations $\theta=\{\theta_{k}\}_{k=1}^{K}$. The objective of the design problem is to strategically select points at which to observe the total signal to infer the locations of the source effectively. More implementation details can be found in Section C.2, and Figure 3 is in Appendix D.
Relative performance of BAD and DAD Like in the toy example, Figure 3 illustrates that covariate shift does not impact generalization performance when the model is well-specified. In the well-specified model, Random and BAD result in similar generalization errors. This may be due to poor estimation of the expected information gain (EIG) in high-dimensional spaces.
Interestingly, under model misspecification, we observe that the generalization error of both BAD and DAD slightly increases (Figure 3(b)) while the degree of covariate shift these methods induce decreases (Figure 3(a)).222 In Figure 3(a), the line corresponding to Random in the well-specified case is hidden by (exactly overlaps with) the line corresponding to Random in the misspecified case. We speculate that this discrepancy is due to the amplifying properties of the designs selected by these methods. In the presence of model misspecification, BAD induces less covariate shift than DAD before $T=20$ but nevertheless results in a higher generalization error. This implies that covariate shift is not the only factor influencing generalization error, and that DAD may perform better than traditional BAD due to its selection of de-amplifying (and potentially unrepresentative) training data.
Relative performance of BAD-Adj Figure 1(f) shows that our novel acquisition function leads to smaller generalization error than BAD. Figure 1(c) shows that the designs selected on the basis of this robust acquisition function are more representative than those selected by BAD.
6 Conclusion
In this paper, we explore the effect of model misspecification on BOED methods. Unlike previous works, which only explored the representativeness of training samples, we introduce a new term “error (de-)amplification” and characterize the contributions of “(de-)amplifying” samples to robustness to model misspecification.
Limitations and future work A limitation of our work is that the main contributions are supported by insights from Theorem 3.7, which provides an upper bound on generalization performance. The degree to which this reflects actual generalization performance depends on the tightness of these bounds. Assessing the tightness of these bounds is therefore an important direction for future work. Another limitation of our work is not incorporating an explicit measure of amplification risk into our proposed acquisition function. Unlike representativeness, the learner cannot assess the degree to which the training samples are (de-)amplifying since this requires access to the (unknown) best fitting approximation. An obvious extension of our work would be to investigate if a nonparametric model can be used to capture the model’s misspecifications, which could potentially guide the selection of (de-)amplifying designs in a principled and automatic manner.
Appendix A Proof of Proposition 3.2
$R_{\mathrm{test}}(f)$ $:=\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(f(\xi)-f^{\star}(\xi))^{2}\right]$ (11)
$=\mathbb{E}_{\xi\sim d_{\text{test}}}\left(\left(\left[f(\xi)-\bar{f}(\xi)\right]+\left[\bar{f}(\xi)-f^{\star}(\xi)\right]\right)^{2}\right)$
$=\underbrace{\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\bar{f}(\xi)-f^{\star}(\xi))^{2}\right]}_{\text{Misspecification Bias}}+\underbrace{\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(f(\xi)-\bar{f}(\xi))^{2}\right]}_{\text{Estimation Bias}}$
$\quad+\underbrace{2\,\mathbb{E}_{\xi\sim d_{\text{test}}}\left[(\bar{f}(\xi)-f^{\star}(\xi))(f(\xi)-\bar{f}(\xi))\right]}_{\text{Error (De-)amplification}}.$
Appendix B Proof of Theorem 3.7
Theorem: with probability at least $1-\delta$, the generalization error under covariate shift satisfies:
$R_{\text{test}}(\hat{f}^{(n)})\leq\mathbb{C}_{\infty}\cdot\left\{\begin{aligned} &\mathbb{B}^{2}+\frac{224y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-2\mathbb{A}(\hat{f}^{(n)}),&\text{if }\mathbb{A}<0\\ &\mathbb{B}^{2}+\frac{128y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-\sqrt{3}\mathbb{A}(\hat{f}^{(n)}),&\text{if }\mathbb{A}\geq 0\end{aligned}\right.$ (12)
Proof of Theorem
We provide the full derivation of the generalization error bound that preserves the error (de-)amplification term. Our analysis, which handles misspecification, is adapted from the proof of Proposition 2.1 in [amortila2024mitigating]. In the proof, we first analyse the empirical error considering the training data distribution $d_{\text{train}}$, and then use the bounded density ratio to analyse the generalization error in the testing data distribution $d_{\text{test}}$.
The goal is to bound the generalization risk
$R(f):=\mathbb{E}_{\xi}[(f(\xi)-f^{\star}(\xi))^{2}],$
for any $f\in\mathcal{F}$. And the empirical risk is defined as:
$\quad\widehat{R}(f):=\frac{1}{n}\sum_{i=1}^{n}\left(f\left(\xi_{i}\right)-y_{i}\right)^{2},$
Observe that conditional on any $\xi$ we have:
$\widehat{R}(f)-\widehat{R}(\bar{f})$ (13)
$=\mathbb{E}[(f\left.(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\mid\xi\right]$
$=\mathbb{E}\left[(f(\xi)-y)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)+f^{\star}(\xi)-y\right)^{2}\mid\xi\right]$
$=\mathbb{E}\left[(f(\xi)-y)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}-2\left(\bar{f}(\xi)-f^{\star}(\xi)\right)\left(f^{\star}(\xi)-y\right)-\left(f^{\star}(\xi)-y\right)^{2}\mid\xi\right]$
$=\mathbb{E}\left[(f(\xi)-y)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}-\left(f^{\star}(\xi)-y\right)^{2}\mid\xi\right]$
$=f(\xi)^{2}-f^{\star}(\xi)^{2}-2\mathbb{E}_{\text{train }}[y\mid\xi]\left(f(\xi)-f^{\star}(\xi)\right)-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}$
$=\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\quad(\text{since}\quad\mathbb{E}_{\text{train }}[y\mid\xi]=f^{\star}(\xi))$
thus
$\mathbb{E}(\widehat{R}(f)-\widehat{R}(\bar{f}))=\mathbb{E}(\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2})=R(f)-R(\bar{f})$
And
$\operatorname{Var}\left[(f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right]$ (14)
$=\mathbb{E}\left[\left((f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right)^{2}\right]-\left(\mathbb{E}\left[\left((f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right)\right]\right)^{2}\quad(\text{since}\quad\operatorname{Var}(X)=\mathbb{E}(X^{2})-(\mathbb{E}X)^{2})$
$\leq\mathbb{E}\left[\left((f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right)^{2}\right]$
$=\mathbb{E}\left[\left(f^{2}(\xi)-2f(\xi)y-\bar{f}^{2}(\xi)+2y\bar{f}(\xi)\right)^{2}\right]$
$=\mathbb{E}\left[(f(\xi)-\bar{f}(\xi))^{2}(f(\xi)+\bar{f}(\xi)-2y)^{2}\right]$
$\leq 16y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-\bar{f}(\xi))^{2}\right]\quad\text{(since }|f(\xi)|,|\bar{f}(\xi)|,|y|\leq y_{\infty}\text{)}$
$=16y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-f^{\star}(\xi)-(\bar{f}(\xi)-f^{\star}(\xi)))^{2}\right]$
$=16y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}+\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]-32y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-f^{\star}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$
$=16y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}+2\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]-32y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-f^{\star}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$
$=16y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}+2\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]$
$\quad-32y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-\bar{f}(\xi)+\bar{f}(\xi)-f^{\star}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$
$=16y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]+32y_{\infty}^{2}\mathbb{E}\left(\bar{f}(\xi)-f^{\star}(\xi)^{2}\right)$
$\quad-32y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]-32y_{\infty}^{2}\mathbb{E}[(\bar{f}(\xi)-f^{\star}(\xi))^{2}]$
$\leq 16y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]-32y_{\infty}^{2}\underbrace{\mathbb{E}\left[(f(\xi)-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]}_{\mathbb{A}(f)}$
$=16y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]-32y_{\infty}^{2}\mathbb{A}(f)$
$=16y_{\infty}^{2}(R(f)-R(\bar{f}))-32y_{\infty}^{2}\mathbb{A}(f)$
Application of Bernstein’s Inequality and Union Bound
From Bernstein’s inequality [shalev2014understanding] and union bound (see Lemma 2.2 in [shalev2014understanding])
Thus, the Bernstein’s inequality becomes:
$\mathbf{P}\left(\mathbf{E}Z^{(f)}-\frac{1}{n}\sum_{i=1}^{n}Z^{(f)}_{i}<\frac{c}{3n}\log(|\mathcal{F}|/\delta)+\sqrt{\frac{2(\operatorname{Var}Z^{(f)})\log(|\mathcal{F}|/\delta)}{n}}\right)\geq 1-\delta$
Setting $Z^{(f)}_{i}=(f(\xi)-y_{i})^{2}-(\bar{f}(\xi)-y_{i})^{2}$, thus $\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{(f)}=\widehat{R}(f)-\widehat{R}(\bar{f});\quad\mathbb{E}Z^{(f)}=R(f)-R(\bar{f})$ , $|Z^{(f)}-EZ^{(f)}|\leq 2\sup|Z^{(f)}|\leq 2\sup(f(\xi)-y_{i})^{2}-(\bar{f}(\xi)-y_{i})^{2}=2\sup(f(\xi)-y_{i}+\bar{f}(\xi)-y_{i})(f(\xi)-\bar{f}(\xi))=2\cdot(4y_{\infty})(2y_{\infty})=c$, thus $c=16y_{\infty}^{2}$
Now, by using Bernstein’s inequality and a union bound over $f\in\mathcal{F}$ that with probability at least $1-\delta$
$\forall f\in\mathcal{F}:R(f)-R(\bar{f})-(\widehat{R}(f)-\widehat{R}(\bar{f}))\leq\sqrt{\frac{\left(16(R(f)-R(\bar{f}))-32\mathbb{A}(f)\right)2y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{n}}+\frac{16y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}.$ (15)
Supppose the
$\widehat{R}\left(\hat{f}^{(n)}\right)-\widehat{R}(\bar{f})\leq 0$ (16)
we have
$R\left(\hat{f}^{(n)}\right)-R(\bar{f})$ $\leq\sqrt{\frac{\left(16\left(R\left(\hat{f}^{(n)}\right)-R(\bar{f})\right)-32\mathbb{A}(\hat{f}^{(n)})\right)2y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{n}}+\frac{16y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}$ (17)
To simplify, make
$A:=R(\hat{f}^{(n)})-R(\bar{f})$ (18)
$C:=y_{\infty}^{2}\frac{\log(|\mathcal{F}|/\delta)}{n}>0$
Why $(16A-32\mathbb{A})\geq 0$
$A$ $:=R(\hat{f}^{(n)})-R(\bar{f})=\mathbb{E}_{\xi}[(\hat{f}^{(n)}(\xi)-f^{\star}(\xi))^{2}]-\mathbb{E}_{\xi}[(\bar{f}(\xi)-f^{\star}(\xi))^{2}]$ (19)
$=\mathbb{E}_{\xi}[(\hat{f}^{(n)}(\xi)-f^{\star}(\xi)+\bar{f}(\xi)-f^{\star}(\xi))(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))]$
Then
$1/2A-\mathbb{A}$ $=1/2\mathbb{E}_{\xi}[(\hat{f}^{(n)}(\xi)-f^{\star}(\xi)+\bar{f}(\xi)-f^{\star}(\xi))(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))]-\mathbb{E}_{\xi}\left[(\hat{f}^{(n)}-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$ (20)
$=\mathbb{E}_{\xi}[(1/2\hat{f}^{(n)}(\xi)-1/2f^{\star}(\xi)+1/2\bar{f}(\xi)-1/2f^{\star}(\xi)-\bar{f}(\xi)+f^{\star}(\xi))(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))]$
$=\mathbb{E}_{\xi}[(1/2\hat{f}^{(n)}(\xi)-1/2\bar{f}(\xi))(\hat{f}^{(n)}(\xi)-\bar{f}(\xi))]\geq 0$
B.1 Suppose that $\mathbb{A}<0$
So the above inequality function equals:
$A$ $\leq\sqrt{(16A-32\mathbb{A})2C}+\frac{16C}{3}$ (21)
$\leq\sqrt{32AC}+\sqrt{-64\mathbb{A}C}+\frac{16C}{3}\quad(\text{since}\sqrt{a+b}<\sqrt{a}+\sqrt{b})$
$=\sqrt{A\cdot 32C}+\sqrt{-2\mathbb{A}\cdot 32C}+\frac{16C}{3}$
$\leq\frac{1}{2}A+16C-\mathbb{A}+16C+\frac{16C}{3}\quad(\sqrt{ab}\leq a/2+b/2)$
$=\frac{1}{2}A-\mathbb{A}+\frac{112C}{3}$
so that
$A\leq-2\mathbb{A}+\frac{224C}{3}$ (22)
So
$R\left(\hat{f}^{(n)}\right)-R(\bar{f})\leq-2\mathbb{A}(\hat{f}^{(n)})+\frac{224y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}$ (23)
Re-arranging and using that ${R}_{\text{train}}(\bar{f})\leq\mathbb{B}^{2}$, thus we have
$R_{\text{train}}\left(\hat{f}^{(n)}\right)\leq\mathbb{B}^{2}-2\mathbb{A}(\hat{f}^{(n)})+\frac{224y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}$ (24)
Finally, Let $\mathbb{C}_{\infty}:=\sup_{\xi\in\Xi}\left|\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\right|$, the upper bound of generalization error can be expressed via the density ratio:
$R_{\text{test}}\left(\hat{f}^{(n)}\right)=\mathbb{E}_{\text{train}}\left[\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\left(\hat{f}^{(n)}(\xi)-f^{*}(\xi)\right)^{2}\right]\leq\mathbb{C}_{\infty}\cdot\left(\mathbb{B}^{2}-2\mathbb{A}(\hat{f}^{(n)})+\frac{224y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}\right)$ (25)
where $\mathbb{A}(\hat{f}^{(n)})=\mathbb{E}_{\text{train}}\left[(\hat{f}^{(n)}-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$
B.2 Suppose that $\mathbb{A}\geq 0$
Although the second solution, based on solving a quadratic equation, yields a tighter upper bound, it involves more complex derivations. In contrast, the first solution offers a simpler and more interpretable decomposition, while preserving the same theoretical insights into the behavior of the error (de-)amplification term $\mathbb{A}(\hat{f}^{(n)})$. Therefore, we adopt the first formulation in the main text and provide the alternative quadratic-based bound in the appendix for completeness.
B.2.1 Solution 1: First order based upper bound derivation
Suppose that $\mathbb{A}\geq 0$ and $A>0$. We use the first-order upper bound for concave functions [boyd2004convex]:
$\sqrt{a+b}\leq\sqrt{a}+\frac{b}{2\sqrt{a}}\quad\text{for }a>0,\ a+b\geq 0.$
Since $(16A-32\mathbb{A})\geq 0$, thus $0\leq\mathbb{A}\leq 1/2A$
$A$ $\leq\sqrt{(16A-32\mathbb{A})2C}+\frac{16C}{3}$ (26)
$\leq\sqrt{32AC}-\frac{64\mathbb{A}C}{2\sqrt{32AC}}+\frac{16C}{3}$
$=\sqrt{A\cdot 32C}-\frac{4\sqrt{2}\mathbb{A}\sqrt{C}}{\sqrt{A}}+\frac{16C}{3}$
$\leq\frac{1}{2}A+16C-\frac{4\sqrt{2}\mathbb{A}\sqrt{C}}{\sqrt{A}}+\frac{16C}{3}\quad(\text{since}\sqrt{ab}\leq a/2+b/2)$
$=\frac{1}{2}A-\frac{4\sqrt{2}\mathbb{A}\sqrt{C}}{\sqrt{A}}+\frac{64C}{3}$
$\leq\frac{1}{2}A-\frac{4\sqrt{2}\mathbb{A}\sqrt{C}}{\sqrt{A_{max}}}+\frac{64C}{3}$
so that
$A$ $\leq-\frac{8\sqrt{2}\mathbb{A}\sqrt{C}}{\sqrt{A_{max}}}+\frac{128C}{3}$ (27)
$=-\frac{8\sqrt{2}\mathbb{A}\sqrt{C}}{\sqrt{\frac{128C}{3}}}+\frac{128C}{3}\quad(\text{make}\mathbb{A}=0,\text{thus}A_{max}<\frac{128C}{3})$
$=\frac{128C}{3}-\sqrt{3}\mathbb{A}$
thus:
$R\left(\hat{f}^{(n)}\right)-R(\bar{f})$ $\leq\frac{128y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-\sqrt{3}\mathbb{A}$ (28)
Re-arranging and using that ${R}_{\text{train}}(\bar{f})\leq\mathbb{B}^{2}$, thus we have
$R_{train}\left(\hat{f}^{(n)}\right)$ $\leq\mathbb{B}^{2}+\frac{128y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-\sqrt{3}\mathbb{A}$ (29)
Finally, Let $\mathbb{C}_{\infty}:=\sup_{\xi\in\Xi}\left|\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\right|$, the upper bound of generalization error can be expressed via the density ratio:
$R_{test}\left(\hat{f}^{(n)}\right)$ $\leq\mathbb{C}_{\infty}\cdot\left(\mathbb{B}^{2}+\frac{128y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}-\sqrt{3}\mathbb{A})\right)$ (30)
where $\mathbb{A}(\hat{f}^{(n)})=\mathbb{E}_{\text{train}}\left[(\hat{f}^{(n)}-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$
B.2.2 Solution 2: Quadratic-based upper bound derivation
Or solving the quadratic inequality $A\leq\sqrt{(16A-32\mathbb{A})2C}+\frac{16C}{3}$, thus
$(A-\frac{16C}{3})^{2}\leq 32AC-64\mathbb{A}C$ (31)
$\Rightarrow A^{2}-\frac{32AC}{3}+(\frac{16C}{3})^{2}\leq 32AC-64\mathbb{A}C$
$\Rightarrow A^{2}-\frac{128AC}{3}+(\frac{16C}{3})^{2}+64\mathbb{A}C\leq 0$
thus,
$A\leq\frac{64C}{3}+\sqrt{\frac{1280C^{2}}{3}-64\mathbb{A}C}$ (32)
[…]
thus:
$R\left(\hat{f}^{(n)}\right)-R(\bar{f})$ $\leq\frac{64y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}+\sqrt{\frac{1280y_{\infty}^{4}[\log(|\mathcal{F}|/\delta)]^{2}}{3n^{2}}-\frac{64\mathbb{A}y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{n}}$ (33)
Re-arranging and using that ${R}_{\text{train}}(\bar{f})\leq\mathbb{B}^{2}$, thus we have
$R_{train}\left(\hat{f}^{(n)}\right)$ $\leq\mathbb{B}^{2}+\frac{64y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}+\sqrt{\frac{1280y_{\infty}^{4}[\log(|\mathcal{F}|/\delta)]^{2}}{3n^{2}}-\frac{64\mathbb{A}y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{n}}$ (34)
Finally, Let $\mathbb{C}_{\infty}:=\sup_{\xi\in\Xi}\left|\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\right|$, the upper bound of generalization error can be expressed via the density ratio:
$R_{test}\left(\hat{f}^{(n)}\right)$ $\leq\mathbb{C}_{\infty}\cdot\left(\mathbb{B}^{2}+\frac{64y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}+\sqrt{\frac{1280y_{\infty}^{4}[\log(|\mathcal{F}|/\delta)]^{2}}{3n^{2}}-\frac{64\mathbb{A}y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{n}})\right)$ (35)
where $\mathbb{A}(\hat{f}^{(n)})=\mathbb{E}_{\text{train}}\left[(\hat{f}^{(n)}-\bar{f}(\xi))(\bar{f}(\xi)-f^{\star}(\xi))\right]$
B.3 loosen the $\mathbb{A}$
We reproduce this derivation for completeness, following the approach in [amortila2024mitigating], although it is not directly used in our main results.
In the inequality 14, we keep the interaction term $\mathbb{A}$. and if we loose the $\mathbb{A}$ by using the AM-GM inequality $\left((a+b)^{2}\leq 2a^{2}+2b^{2}\right)$, the above equation becomes :
$\operatorname{Var}\left[(f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right]$ (36)
$=\mathbb{E}\left[\left((f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right)^{2}\right]-\left(\mathbb{E}\left[\left((f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right)\right]\right)^{2}$
$\leq\mathbb{E}\left[\left((f(\xi)-y)^{2}-(\bar{f}(\xi)-y)^{2}\right)^{2}\right]$
$=\mathbb{E}\left[\left(f^{2}(\xi)-2f(\xi)y-\bar{f}^{2}(\xi)+2y\bar{f}(\xi)\right)^{2}\right]$
$=\mathbb{E}\left[(f(\xi)-\bar{f}(\xi))^{2}(f(\xi)+\bar{f}(\xi)-2y)^{2}\right]$
$\leq 16y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-\bar{f}(\xi))^{2}\right]\quad\text{(since }|f(\xi)|,|\bar{f}(\xi)|,|y|\leq y_{\infty}\text{)}$
$=16y_{\infty}^{2}\mathbb{E}\left[(f(\xi)-f^{\star}(\xi)+f^{\star}(\xi)-\bar{f}(\xi))^{2}\right]$
$\leq 32y_{\infty}^{2}\mathbb{E}\left[((f(\xi)-f^{\star}(\xi))^{2}+(\bar{f}(\xi)-f^{\star}(\xi))^{2})\right]\quad\text{(since }(a+b)^{2}\leq 2a^{2}+2b^{2})$
$=32y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}+2\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]$
$\leq 32y_{\infty}^{2}\mathbb{E}\left[\left(f(\xi)-f^{\star}(\xi)\right)^{2}-\left(\bar{f}(\xi)-f^{\star}(\xi)\right)^{2}\right]+64y_{\infty}^{2}\mathbb{B}^{2}$
$=32y_{\infty}^{2}(R(f)-R(\bar{f}))+64y_{\infty}^{2}\mathbb{B}^{2}$
Using the same logic in the above derivation, based on the LABEL:equ:loosA, we have:
$A$ $\leq\sqrt{(32A+64B)2C}+\frac{16C}{3}$ (37)
$\leq\sqrt{64AC}+\sqrt{128BC}+\frac{16C}{3}\quad(\text{since}\sqrt{a+b}<\sqrt{a}+\sqrt{b})$
$=\sqrt{A\cdot 64C}+\sqrt{2B\cdot 64C}+\frac{16C}{3}$
$\leq\frac{1}{2}A+32C+B+32C+\frac{16C}{3}\quad(\text{since}\sqrt{ab}\leq a/2+b/2)$
$=\frac{1}{2}A+B+\frac{208C}{3}$
so that
$A\leq 2B+\frac{416C}{3}$ (38)
So
$R\left(\hat{f}^{(n)}\right)-R(\bar{f})\leq 2\mathbb{B}^{2}+\frac{416y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}$ (39)
Re-arranging and using that ${R}_{\text{train}}(\bar{f})\leq\mathbb{B}^{2}$, thus we have
$R_{\text{train}}\left(\hat{f}^{(n)}\right)\leq 3\mathbb{B}^{2}+\frac{416y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}$ (40)
Let $\mathbb{C}_{\infty}:=\sup_{\xi\in\Xi}\left|\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\right|$, the upper bound of generalization error can be expressed via the density ratio, :
$R_{\text{test}}\left(\hat{f}^{(n)}\right)=\mathbb{E}_{\text{train}}\left[\frac{d_{\text{test}}(\xi)}{d_{\text{train}}(\xi)}\left(\hat{f}^{(n)}(\xi)-f^{*}(\xi)\right)^{2}\right]\leq\mathbb{C}_{\infty}\cdot\left(3\mathbb{B}^{2}+\frac{416y_{\infty}^{2}\log(|\mathcal{F}|/\delta)}{3n}\right)$ (41)
Appendix C Experiment details
Our experiments were implemented using PyTorch [paszke2019pytorch] and Pyro [bingham2019pyro]. The implementation of the DAD method is based on the official code provided by [foster2021deep]. All experiments were conducted on a shared computing cluster using NVIDIA A100 GPUs. Each job was allocated one A100 GPU, 8 CPU cores, and 32GB of RAM.
Measuring the degree of covariate shift To measure the distance between the distributions or datasets, we use the Maximum mean discrepancy (MMD). A number of advantages of this distance are put forward in the literature: 1) it is more robust to outliers than other discrepancy measurements (like KL divergence) [gretton2012kernel]; 2) it can be approximated on the basis of differently-sized samples from the distributions being compared [gretton2012kernel]; 3) the measurement is robust to repeated samples (unlike the Wasserstein distance) [viehmann2021partial]; 4) the measurement can be computed efficiently using samples [bharti2023optimally, huang2023learning].
C.1 Toy example
Both well-specified and misspecified models were run across 20 runs. For each run, the design is adaptively selected, and the total number of designs is $T=10$.
Well-specified case
  • DGP: The DGP is a degree-two polynomial regression model, $y=1+2x-0.5x^{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,0.1)$
  • well-specified model: $f(x)=\beta_{0}+\beta_{1}x+\beta_{2}x^{2}$ So the feature functions are $\phi(x)=[1,x,x^{2}]^{\top}$
  • Test distribution (arbitrary $\mathbf{\xi}^{test}$): $\mathbf{\xi}^{test}\sim\mathcal{U}(-4,4)$
Misspecified case
  • DGP: The DGP is a degree-two polynomial regression model, $y=1+2x-0.5x^{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,0.1)$
  • misspecified model: $f(x)=\beta_{0}+\beta_{1}x$ So the feature functions are $\phi(x)=[1,x]^{\top}$
  • Test distribution (arbitrary $\mathbf{\xi}^{test}$): $\mathbf{\xi}^{test}\sim\mathcal{U}(-4,4)$
C.2 Source localization experiments
Implementation details We take $T=30$ iteration steps for selecting training samples $\mathbf{\xi}^{train}$ and 200 test samples for the observed $f^{\star}(\mathbf{\xi}^{test})$ and predicted output $f(\mathbf{\xi}^{test})$. We set $K=2$ sources and limit the range of design to $\mathbf{\xi}^{train}\in[-4,4]$. We use 200 samples drawn from the uniform distribution over $\mathbf{\xi}^{test}\in[-4,4]$ as the test distribution to estimate the MMD.
In this experiment, we have $K$ sources with unknown location parameter is $\theta=\{\theta_{k}\}_{k=1}^{K}$. We assume the number of sources $K$ is known.
Acoustic energy attenuation model
The total intensity, a superposition of the individual ones, at location $\xi$ from $K$ sources is considered in the following equation [foster2021deep, ivanova2024step]:
$\mu(\theta,\xi)=b+\sum_{k=1}^{K}\frac{\alpha_{k}}{m+\left\|\theta_{k}-\xi\right\|^{2}}$ (42)
The prior distribution of each location parameter $\theta_{k}$ is a normal distribution, and the observation noise is Gaussian noise. Therefore, the prior and likelihood are in the following:
$\theta_{k}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}N\left(0_{d},I_{d}\right),\log y\mid\theta,\xi\sim N(\log\mu(\theta,\xi),\sigma)$
Well-specfied case
The hyperparameters of the true DGP used in our well-specification experiments can be found in the table below:
[…]
The model hyperparameters used in our experiments are the same in the above table.
Misspecfied case
The hyperparameters of the true DGP used in our mis-specification experiments can be found in the table below:
[…]
The model hyperparameters used in our experiments are the same as the table in the well-specified case.
Trained DAD network
Both the misspecified and well-specified cases use the same DAD network architecture. The encoder and decoder are both simple feed-forward neural networks with a single hidden layer. Detailed configurations are provided in the following table. For the encoder
[…]
and for the emitter
[…]
Full details of the training are given in the following table.
[…]
Appendix D Experimental results
Figure 3: Comparison of different design strategies (Random, BAD, DAD) under well-specified and misspecified models in source location experiments. Left: MMD distance across methods; higher values indicate a greater degree of covariate shift. Right: Generalization error across methods.